OpenStudy (anonymous):

what is (3,5), (x,2); undefined slope equal?

OpenStudy (abdullahm):

If it has an undefined slope, that means the denominator is 0. So, do you know the slope formula? @jadeisgold7987

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (abdullahm):

Correct, and if it has 0 in the denominator, that means \(\ x_2 - x_1 =0\) And we already know one of the x-values, so can't we figure out the second x-value using that?

OpenStudy (anonymous):

uhhh i think so.....

OpenStudy (abdullahm):

So, one x-value is 3. \(3-x_2=0\) Can you solve for the second x-value?

OpenStudy (anonymous):

its 3

OpenStudy (abdullahm):

Exactly, so x = 3. That's your answer! :)

OpenStudy (anonymous):

yay!!!!!! thanks!

OpenStudy (abdullahm):

No problem!

OpenStudy (anonymous):

ok but wat if the equatio was this\[\frac{ -3 }{ x-3 }\]